Find the solution to each system by the substitution method. 10x + 3y =8 and 2x + y = 2
Answer by drk(1908)

Here are the equations:
(i) 10x + 3y = 8
(ii) 2x + y = 2
Solve (ii) for y. It is easiest.
(i) 10x + 3y = 8
(iii) y = -2x + 2
substitute (iii) into (i) to get
10x + 3(-2x+2) = 8
10x - 6x + 6 = 8
4X = 2
X = 1/2
find Y.
Y = 1.
(1/2, 1)